OpenStudy (anonymous):

Find derivative y=1/square root of x^2-7x+3

OpenStudy (anonymous):

Are you still here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Are you pretty familiar with negative exponents, techdodo?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Fractional exponents, too?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x^2-7 x+3}}\] \[\left(x^2-7 x+3\right)^{\frac{-1}{2}}\] You will have to use chain rule first take derivative of outer function, don't mind what inside. \[\frac{-1}{2}\left(\text{ }x^2-7 x+3\text{ }\right)^{\frac{-3}{2}}\] Now take inner derivative 2x-7 Multiply inner and outer derivative \[\frac{-1}{2}\left(\text{ }x^2-7 x+3\text{ }\right)^{\frac{-3}{2}}(2x-7)\]

OpenStudy (anonymous):

Excellent! Let's do this one step at a time, then: y=1/square root of x^2-7x+3 Can you show me what this would look like in pure exponential form?

OpenStudy (anonymous):

Still here, dodo?

OpenStudy (anonymous):

Ah, oh well. Imran showed you the exact steps I was going to walk you through, so just follow those carefully.

OpenStudy (anonymous):

yea was just having trouble with the radical

OpenStudy (anonymous):

but i forgot bout bringing it up and making the exponent negative

OpenStudy (anonymous):

O ok. Well let's do it one step at a time. How do you transform the radical into an exponent?

OpenStudy (anonymous):

thank you though

OpenStudy (anonymous):

O ok. Glad to know you knew it. :D

OpenStudy (anonymous):

(x^2-7x+3)^1/2

OpenStudy (anonymous):

Yup :D And then you just do the negative, and you're all set to chain rule. Great work. :D

OpenStudy (anonymous):

just the exponent would be negative when it moves up right?

OpenStudy (anonymous):

Correct. Because 1/(anything) means (anything)^-1

OpenStudy (anonymous):

Likewise, 1/((anything)^2) would equal (anything)^-2